# Benchmarking Algorithms for Federated Domain Generalization

### v1

@@ 2.1 Problem Background and Setup
Domain separation ($\lambda=0$) FL with domain separation Federated DG with domain separation
Overview of Federated DG methods
In this benchmark study, we explore three categories of Federated DG methods: DG methods originally designed for the centralized setting, FL methods specifically tailored to handle client heterogeneity, and methods specifically designed for Federated DG. To provide a comprehensive evaluation, we assess the performance of several representative methods from each of these categories and compare to vanilla FedAvg (McMahan et al., 2017) with ERM loss, where any heterogeneity is simply ignored. These methods are selected based on their prominence and potential effectiveness for Federated DG. To ensure a diverse range of evaluation scenarios, we conduct experiments on various datasets and under various heterogeneity settings. Our results provide insights into their relative performance and suitability for different Federated DG contexts. This benchmark study aims to offer a broad overview of the current popular methods, enabling researchers and practitioners to make informed decisions when tackling Federated DG.
2.2 Centralized DG methods adapted to the FL setting
The first natural choice is directly migrating the DG methods from the centralized regime. To adapt those methods, we simply run the centralized DG method at each client locally with their own local dataset (see subsection 3.2 for how the local datasets are created), and then compute an average of model parameters at each communication round (see next paragraph). This approach is straightforward for the homogeneous ($\lambda=1$) and heterogeneous ($\lambda=0.1$) settings where each client has data from all training domains—albeit quite imbalanced for $\lambda=0.1$. This can be seen as biased updates at each client based on biased local data. In the domain separation case $\lambda=0$, (i.e., if all clients only have one primary domain, i.e., $\forall k,|P_{k}|=1$), this simple approach cannot be applied because centralized DG methods require data from at least two domains. In fact, these centralized DG methods degenerate to vanilla FedAvg if there is only one domain per client. Extending these methods to the case where all clients only have one domain without violating the FL constraints is an interesting direction for future work.
A predominant and effective centralized DG approach is through representation learning, including domain-invariant representation learning by learning domain-invariant features via either kernel methods (Muandet et al., 2013; Ghifary et al., 2016) or invariant risk minimization (Arjovsky et al., 2019; Krueger et al., 2020), and domain adversarial neural networks Sun and Saenko (2016), and Li et al. (2018b) which explicitly tries to align the feature representation distribution. Besides invariant representation, there are other methods the general learning strategy to promote the generalization ability gradient operation Shi et al. (2021); Rame et al. (2022), which tries to learn generalized representations by directly operating on gradients. Other approaches include distributionally robust optimization Sagawa et al. (2019), which learns the worst-case distribution scenario of training domains; and meta-learning Finn et al. (2017); Li et al. (2018a), which is based on the learning-to-learn mechanism to learn general knowledge by constructing meta-learning tasks to simulate domain shift. We selected IRM (Arjovsky et al., 2019), Fish (Shi et al., 2021), MMD (Li et al., 2018b), Mixup (Zhang et al., 2017), DeepCoral (Sun and Saenko, 2016), and GroupDRO (Sagawa et al., 2019) from this category for their representative mechanisms.
2.3 FL methods tackling client heterogeneity
Another line of research in FL aims to guarantee convergence even under client heterogeneity, but these FL-based methods still assume the train and test datasets do not shift (i.e., they do not explicitly tackle train-test heterogeneity of the domain generalization task). The empirical observation of the statistical challenge in federated learning when local data is non-IID was first made by Zhao et al. (2018). Several subsequent works have analyzed client heterogeneity by assuming bounded gradients (Yu et al., 2019; Basu et al., 2019; Wang et al., 2019; Li et al., 2019) or bounded gradient dissimilarity (Li et al., 2020), and additionally assuming bounded Hessian dissimilarity (Karimireddy et al., 2020; Khaled et al., 2020; Liang et al., 2019). From this category, we selected FedProx (Li et al., 2020), which addresses statistical heterogeneity by adding a proximal term to the local subproblem, constraining local updates to be closer to the initial global model. Scaffold (Karimireddy et al., 2020) utilizes variance reduction to account for client heterogeneity.
2.4 FL methods designed for Federated DG
Limited research has focused explicitly on solving the Federated DG by design. FedDG Liu et al. (2021) introduced a specific FL paradigm for medical image classification, which involves sharing the amplitude spectrum of images among local clients, violating the privacy protocol. Another approach, FedADG (Zhang et al., 2021), utilizes a generative adversarial network (GAN) framework in FL, where each client contains four models: a featurizer, a classifier, a generator, and a discriminator. FedADG first trains the featurizer and classifier using empirical loss and then trains the generator and discriminator using the GAN approach. However, this method requires training four models simultaneously and tuning numerous hyperparameters, making convergence challenging. A novel aggregation method called Federated Gradient Masking Averaging (FedGMA) (Tenison et al., 2022) aims to enhance generalization across clients and the global model. FedGMA prioritizes gradient components aligned with the dominant direction across clients while assigning less importance to inconsistent components. FedSR (Nguyen et al., 2022) proposes a simple algorithm that uses two locally-computable regularizers for domain generalization. Given the limited literature on solving domain generalization (DG) in the federated learning (FL) setting, we selected all the aforementioned algorithms.
3 Benchmark Methodology
In this study, we aimed to conduct a comprehensive evaluation of the Federated DG task by considering four distinct dimensions of the problem setup. We evaluated a total of $13$ methods, encompassing three different types of approaches.
3.1 Four evaluation dimensions
(1) Domain-based client heterogeneity and (2) number of clients
[…]

### v2

@@ 2.1 Problem Background and Setup
Complete Heterogeneity ($\lambda=0$) FL with complete heterogeneity Federated DG with complete heterogeneity
2.2 Current Data partition methods
In the FL context, the method for partitioning training data is crucial. As of now, there are three primary techniques for dataset partitioning. These methods are all fall in short in some properties we need. See Section 3 for a comprehensive analysis. In this section, we primarily present these methods as a background. Shards partitioning (McMahan et al., 2017) is a deterministic partition method. The datasets are initially ordered according to their labels. Then the datasets are partitioned into $2\times C$ shards, and each client $c\in[C]$ is allocated $2$ shards. Dirichlet Partitioning (Yurochkin et al., 2019; Hsu et al., 2019) is a stochastic partition method where each client hold the same number of samples but the proportion among labels are stochastic following a probability vector sampled from the Dirichlet distribution $\text{Dir}(\alpha p)$, where $p$ represents the prior label distribution on the whole training dataset, and $\alpha$ control the heterogeneity level. As $\alpha\rightarrow 0$, each client predominantly holds samples of a single label. Conversely, as $\alpha\rightarrow\infty$, the distribution of labels for each client becomes identical. Semantic Partitioning (Yuan et al., 2022) tries to create client heterogeneity on the features of data samples. Firstly, the data is processed through a pretrained model. The outputs from this model are then used to fit a Gaussian Mixture Model that features $K$ clusters for each class $Y$. These $KY$ clusters are then merged iteratively using an optimal bipartite for random label pairs.
3 Heterogeneous Partitioning Method
In this section, to thoroughly assess current Federated DG methods, we formulate an optimization problem that an ideal partition method should aim to solve. We then demonstrate that previous partition methods (subsection 2.2) are not feasible solutions. Subsequently, we introduce a new partition method, termed Heterogeneous Partitioning, which is an approximate optimal solution. Generally, our partition method effectively handles partitioning $D$ types of integer-numbered objects into $C$ groups. It’s broadly applicable, suitable for domain adaptation, ensuring fairness in Federated Learning (FL), and managing non-iid FL regimes.
Table 2: Comparison of different partition methods.
$C_{1}$ $C_{2}$
Shards No Yes
Dirichlet Yes No
Semantic No Yes
Ours Yes Yes
Assume we have a set of all samples $\mathcal{S}=\{(x_{i},d_{i})\}_{i=1}^{n}$ and let $\mathcal{S}_{c}$ denote the set of samples assigned to the $c$-th client. For each client $c\in[C],$ define its domains as the set of domains from which $c$ has at least one sample., i.e., $\mathcal{D}_{c}=\{d\in[D]:\exists(x_{i},d_{i})\in\mathcal{S}_{c},d_{i}=d\}$. Denote $\mathcal{P}$ as a partition method.
For Federated DG, it is common that clients have datasets collected from disjoint domains Koh et al. (2021). Therefore, it’s crucial that $\mathcal{P}$ can generate the strongest possible heterogeneity given the datasets and clients to reflect those common scenarios. This requirement can be encoded as either a single domain per client or no intersection of client domains, depending on the numbers of clients and domains (see constraint $C_{1}$). Furthermore, in distributed setting, to ensure a reasonable comparison with respect to domain generalization in the centralized setting, the total datasets received over the distributed system should match that of the centralized setting; and due to the privacy protocol, clients should not share the datasets. Thus, we need to preclude the partitions $\mathcal{P}$ where some data from $\mathcal{S}$ is not used by any client or clients either have duplicated data, which translate to constraint $C_{2}$. Thus, the feasible region of partition $\mathcal{P}$ is defined as $\mathcal{P}\in C_{1}\cap C_{2}.$
Complete Heterogeneity (constraint $C_{1}$): $\mathcal{P}$ can generate the possibly strongest domain heterogeneity given the datasets and clients, i.e., $\mathcal{P}\in C_{1}$ where $C_{1}$ is defined as
$C_{1}\triangleq\left\{\mathcal{P}\,\bigg{|}\,\text{ if $C>D,$ then $|\mathcal{D}_{c}|=1,\forall c\,;$ if $C\leq D,$ then $|\mathcal{D}_{c}\cap\mathcal{D}_{c^{\prime}}|=0,\forall c\neq c^{\prime}.$}\right\}$ (2)
True Partition (constraint $C_{2}$): $\mathcal{P}$ must produce true set partitions $\{\mathcal{S}_{c}\},c\in[C]$. A true set partition is a non-empty, exhaustive, and pairwise disjoint family of sets, i.e., $\mathcal{P}\in C_{2}$ where $C_{2}$ is defined as
$C_{2}\triangleq\left\{\mathcal{P}\,\bigg{|}\,\text{ $\mathcal{S}_{c}\neq\emptyset,\forall c$, $\bigcup_{c=1}^{C}\mathcal{S}_{c}=\mathcal{S}$, and $\mathcal{S}_{c}\cap\mathcal{S}_{c^{\prime}}=\emptyset,\forall c\neq c^{\prime}.$}\right\}.$ (3)
[…]
$\widehat{\mathcal{P}}\in\operatorname*{arg\,min}\limits_{\mathcal{P}\in C_{1}\cap C_{2}}\frac{1}{C-1}\sum_{c=1}^{C}\|n_{c}-\bar{n}\|_{2}^{2},$ (4)
where $n_{c}=|\mathcal{S}_{c}|,$ $\bar{n}=\frac{1}{C}\sum_{c}n_{c}$. Notice that in many scenarios, it is not always able to achieve 00 loss, because $n_{c}\equiv\bar{n}$ may not be feasible for constraints $1$ and $2$ For example, consider the case of complete heterogeneity with $C=D$ but imbalance of domains, it is impossible to achieve balance $n_{c}\equiv\bar{n}$ and complete hetereogeneity $|\mathcal{D}_{c}|=1.$ Most previous partitions forces $n_{c}\equiv\bar{n}$ thus violating the constraints. In particular, Dirichlet Partitioning violates $C_{2}.$ Shard and semantic violate $C_{1}.$ We compare different partition methods in Table 2. See subsection A.2 for a detailed discussion on them.
We construct Heterogeneous Partitioning $\{\mathcal{P}_{\lambda}\},\lambda\in[0,1],$ to solve Eqn. 4. It is parameterized by $\lambda,$ where it can generate Complete Hetereogeneity ($\lambda=0$) and always satisfy True Partition (for all $\lambda\in[0,1]$) by construction; further it aims to achieve the best possible balance given these constraints. Heterogeneous Partitioning allows samples allocation from multiple domains across an arbitrary number of clients $C$ while controlling the amount of domain-based client heterogeneity via parameter $\lambda$, ranging from homogeneity to complete heterogeneity. In addition, when $\lambda=0,$ Heterogeneous Partitioning is optimal for Eqn. 4 when there’s more clients than domains $C>D,$ Further, when $C\leq D,$ Eqn. 4 is NP-hard problem proposed by Graham (1969) and Heterogeneous Partitioning is a linear time greedy approximation. The proof is deferred to subsection A.1. Heterogeneous Partitioning contains the following two steps, see Algorithm 1 for pseudo code.
Step 1: Constructing complete heterogeneity. We explicitly construct $\mathcal{P}_{0}$ to satisfy $C_{1}$ while greedily trying to balance the variance. The key idea to create the complete heterogeneity is to partition one domain among the fewest possible clients.
We adopt the following approaches depending on the size of the network $C$ and total domains $D$.
Case I: If $C\leq D$, the domains are sorted in a descending order according to number of sample size in each domain, and are iteratively assigned to the client $c^{*}$ which currently has the smallest number of training samples, that is
$c^{*}\triangleq\operatorname*{arg\,min}\limits_{c\in[C]}\sum_{d^{\prime}\in\mathcal{D}_{c}}n_{d^{\prime}},$ (5)
where $n_{d}$ denotes the total sample size of domain $d.$ In this case, $\mathcal{P}_{0}$ ensures that no client shares domains with the others, i.e., $|D_{c}\cap D_{c}^{\prime}|=0,$ but attempts to balance the number of training samples between clients.
Case II: If $C>D$, we first assign all the domains one by one to the first $D$ clients; then, starting from client $c=D+1$, we divide the currently largest average domain $d^{*}$ between the new client and the original clients associated with it. That is, for all $c\in\{D+1,D+2,\dots,C\},$ we assign $d^{*}$ to $c,$ and reassign the same sample sizes for those $c^{\prime}\in\{1,\dots,c\}$ which share the domain $d^{*}$ as
$d^{*}\in\operatorname*{arg\,max}\limits_{d\in[D]}\frac{n_{d}}{{\textstyle\sum_{c^{\prime}=1}^{C}}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]},\quad\frac{n_{d^{*}}}{{\textstyle\sum_{c^{\prime}=1}^{C}}\mathbbm{1}[d^{*}\in\mathcal{D}_{c^{\prime}}]},$ (6)
where rounding to integers is carefully handled when not perfectly divisible. Notice that in this case, some clients may share one domain, but no client holds two domains simultaneously, that is $|D_{c}|=1,$ for $c\in[C].$ In this case, $\mathcal{P}_{0}$ attempts to balance the number of samples across clients as much as possible while partitioning one domain among the fewest possible clients.
Step 2: Computing domain sample counts for each domain with interpolation. We define $\mathcal{P}_{\lambda}$ through the sample counts for each client $c\in[C]$ per domain $d\in[D]$, denoted as $n_{d,c}(\lambda)$ based on the balancing parameter $\lambda\in[0,1]:$
$\mathcal{P}_{\lambda}:\quad n_{d,c}(\lambda)=\underbrace{\lambda\frac{n_{d}}{C}}_{\text{homogeneous}}+\underbrace{(1-\lambda)\frac{\mathbbm{1}[d\in\mathcal{D}_{c}]}{\sum_{c^{\prime}=1}^{C}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]}n_{d}}_{\text{complete heterogeneous}}\,,$ (7)
where $\mathcal{D}_{c},c\in[C]$ are constructed by $\mathcal{P}_{0}$ Step $1$ This is simply a convex combination between homogeneous clients ($\lambda=1$) and extreme heterogeneous ($\lambda=0$). A smaller value of $\lambda$ corresponds to a more heterogeneous case. Given the number of samples per client $i\in[C]$ per domain $d\in[D]$, we simply sample $n_{d,c}$ without replacement from the corresponding domain datasets and build up the client datasets. Step $2$ ensures our partition $\mathcal{P}_{\lambda}$ satisfy the constraint $C_{2},$ because all samples are selected (as proven by $\sum_{c}n_{d,c}=n_{d})$ and there is no overlap (sample without replacement), and no client has zero samples given $|D_{c}|>0$ in the $\lambda=0,1$ cases.
4 Benchmark Methodology and Evaluations
In this section, we aim to conduct a comprehensive evaluation of the Federated DG task by considering four distinct dimensions of the problem setup equipped by Heterogeneous Partitioning. The four dimensions are 1) dataset type; 2) data difficulty; 3) domain-based client heterogeneity; and 4) the number of clients. This section is organized as follows: we introduce our benchmark datasets subsection 4.1 covering the first two dimensions, next we introduce $14$ methods we included in our evaluation from three different approaches subsection 4.2, then we introduce our benchmark setting and the evaluation results on all selected methods in subsection 4.3 brought the domain-based client heterogeneity and number of clients into consideration.
4.1 Dataset Type and Dataset Difficulty Metrics
While most Federated DG work focuses on standard image-based datasets, we evaluate methods over a diverse selection of datasets. These datasets encompass multi-domain datasets from simpler, pseudo-realistic ones to the considerably more challenging realistic ones. Specifically, our study includes five image datasets and two text datasets. Additionally, within these $7$ datasets, we include one subpopulation shift within image datasets (CelebA) and another within text datasets (Civilcomments). Furthermore, our dataset selections span a diverse range of subjects including general objects, wild camera traps, medical images, human faces, online comments, and programming codes. We also introduce dataset difficulty metrics to measure the empirical challenges of each dataset in the Federated DG task.
Dataset Difficulty Metrics. To ensure a comprehensive evaluation of current methods across different difficulty levels, we have curated a range of datasets with varying complexities. We define two dataset metrics to measure the dataset difficulty with respect to the DG task and with respect to the FL context using the baseline objective ERM. For DG difficulty, we compute $R_{\text{DG}},$ the ratio of the ERM performance without and with samples from the test domain (i.e., the later is able to “cheat” by seeing part of test domain samples during training). For FL difficulty, we attempt to isolate the FL effect by computing $R_{\text{FL}}(\lambda),$ the ratio of ERM-based FedAvg $\lambda$ client heterogeneity over centralized ERM on in-domain test samples. These dataset difficulty metrics can be formalized as follows, for all $\lambda\in[0,1]$
$R_{\text{DG}}\triangleq\frac{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}^{\prime}_{\text{DG-test}})}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}}\cup{\mathcal{S}^{\prime\prime}_{\text{DG-test}}},\mathcal{S}^{\prime}_{\text{DG-test}})},R_{\text{FL}}(\lambda)\triangleq\frac{\texttt{FedAvg\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}};\lambda)}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}})},$
where ERM_Perf is the performance of ERM using the first argument as training and the second for test, FedAvg_Perf is similar but with the client heterogeneity parameter $\lambda$, $\mathcal{S}_{\text{DG-train}}$ denotes samples from the training domains $\mathcal{D}_{\text{train}},$ $\mathcal{S}_{\text{DG-test}}$ denotes samples from the test domains $\mathcal{D}_{\text{test}},$ and $\mathcal{S}^{\prime}_{\text{DG-test}}$ and $\mathcal{S}^{\prime\prime}_{\text{DG-test}}$ are $20\%,80\%$ partition respectively of $\mathcal{S}_{\text{DG-test}}$. For $R_{\text{FL}}(\lambda)$, we use $\mathcal{S}_{\text{IN-test}}$ (test samples from the training domains) instead of $\mathcal{S}_{\text{DG-test}}$ to isolate the FL effect from the DG effect. We apply these metrics in our $7$ selected datasets and include the summary table in Table 6. Smaller $R_{\text{DG}}$ and $R_{\text{FL}}(\lambda)$ indicate a more challenging dataset. For instance, FEMNIST has $R_{\text{DG}}=1$ indicates the lack of domain heterogeneity. To contrast, IwildCam and CivilComments have small $R_{\text{FL}}$ showing the challenges from the FL side.
4.2 Benchmark Methods
In this benchmark study, we explore three categories of Federated DG methods. The first category is centralized DG methods adapted into FL regimes. The second category is FL methods tackling client heterogeneity and the third category is Federated DG methods. Please see Appendix B for a detailed discussion on current available methods to solve Federated DG. To provide a comprehensive evaluation, we assess the performance of several representative methods from each of these categories. specifically, we choose IRM (Arjovsky et al., 2019), Fish (Shi et al., 2021), Mixup (Zhang et al., 2017), MMD (Gretton et al., 2006), Coral (Sun and Saenko, 2016), GroupDRO (Sagawa et al., 2019) and adapted them into FL regime by applying them in every local clients and keeping the aggregation identical to FedAvg, i.e., weighted averaging aggregation. All those methods collapse to FedAdam if locally there is only one domain available. We consider $4$ methods. FedDG (Liu et al., 2021), FedADG (Zhang et al., 2021), FedSR (Nguyen et al., 2022) and FedGMA (Tenison et al., 2022a), which are naturally designed to solve Federated DG. We also consider FedProx (Li et al., 2020) and Scaffold (Karimireddy et al., 2020) from the Federated methods tackling client heterogeneity approach. All those methods will be compared to two baselines. ERM objective with Adam optimizer (Kingma and Ba, 2014) and its FL counterpart FedAdam (Reddi et al., 2020), which is a variant of FedAvg (McMahan et al., 2017).
4.3 Main Results
In this section, we present the performance results of $14$ representative methods, derived from distinct research areas, on $7$ diverse datasets. For each dataset, we fix the total computation and communication rounds for different methods for a fair comparison.
Client number Equipped with our Heterogeneous Partitioning, we can explore various levels of client heterogeneity and relax the assumption that $C=D$ so that we can leverage both pseudo-realistic and real-world datasets and evaluate methods on larger scale of FL clients. In particular, we set the number of clients to $100$
Validation domain In DG task, we cannot access the test domain data. However, we are particularly concerned about the model performance outside the training domains, thus we preserve a small portion of the domains we can access as held-out validation domains, and the held-out validation domains are used for model selection and early stopping. Please see Appendix D for more detail.
After training, we choose the model according to the early-stopping at the communication round which achieves the best held-out-domain performance, and finally we evaluate the performance on the test-domain in Table 3, Table 4 and Table 5. See subsection D.4 for detailed hyperparameters choices. We make the following remarks on the main results from Table 3, Table 4 and Table 5. Because FEMNIST has low DG difficulty, we defer the results on FEMNIST in the Appendix (Table 11).
Remark 4.1.
FedAvg with an ERM objective is a strong baseline, especially on the general DG datasets. We observe FedAvg-ERM outperforms other methods multiple times. It still servers a strong baseline that is challenging to beat across general DG datasets, similar to the centralized case stated in DomainBed (Gulrajani and Lopez-Paz, 2020) and WILDS (Koh et al., 2021). We recommend always including FedAvg as a baseline in all future evaluations.
Remark 4.2.
Most centralized DG methods degrade in the FL setting. For image datasets, the DG methods adapted to the FL setting (except for GroupDRO) show significant degradation in performance compared to their centralized counterparts as can be seen when comparing the $C=1$ column to the $C>1$ columns in Table 3. Further, degradation can be seen in PACS and CelebA when moving from the homogeneous client setting ($\lambda=1$) to the heterogeneous client setting ($\lambda=0.1$).
Remark 4.3.
FL methods tackling client heterogeneity help convergence. Notably, IWildCam and CivilComments datasets bring greater challenges in the model convergence as Table 6. In this scenario, FL methods tackling the client heterogeneity are able to better tackle this challenge while other methods all failed in this context, see Table 3 and Table 5. Given the fact that this is a common challenge in Federated DG, there’s a need for future research to simultaneously enhance the performance with both client heterogeneity and train-test heterogeneity.
Remark 4.4.
Federated DG methods still in need. Upon evaluating current Federated DG methods on larger network scales, they all fail to outperform ERM. Among the four methods we evaluating, FedDG performs the best and could achieve higher accuracy on PACS with $\lambda=0$. However, FedDG requires sharing all local datasets amplitude information to the aggregation server, thus brings privacy concerns. FedADG methods contains tens of hyperparameters, and due to the nature of GAN (Goodfellow et al., 2020), it is challenging to optimize. FedSR and FedGMA also did not outperform ERM objective. It shows a clear demand for further improvement.
Remark 4.5.
Addressing the subpopulation shift could be an initial step. In the centralized setting, all evaluated methods surpassed the performance of ERM on two subpopulation shift datasets: CelebA and CivilComments. However, in a Federated context, only FedDG managed to outpace ERM on CelebA, and yet it couldn’t match the performance of centralized approaches. Further exploration in federated DG is in demand to bridge this gap.
Remark 4.6.
The performance of real-world data significantly degrades as $\lambda$ decreases. This can be seen from IWildCam and Py150 dataset at Table 3 and Table 5. While it is challenging and expensive to run models for IWildCam and Py150, they show the largest differences between methods and demonstrates the real-world challenge of Federated DG. We suggest including IWildCam and Py150 in most future DG evaluations given their unique nature across datasets.
Table 3: Test accuracy on PACS and IWildCam dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. We report the standard deviation among 3 runs. Please see the Table 8 in the appendix for higher precision report.
PACS ($D=2$) IWildCam ($D=243$)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
B
FedAvg-ERM $\bf 0.94\pm 0.01$ $\bf 0.95\pm 0.02$ $\mathbf{0.96\pm 0.02}$ $\mathbf{0.95\pm 0.02}$ $\mathit{0.34\pm 0.00}$ $\mathbf{0.30\pm 0.00}$ $0.25\pm 0.00$ $\mathbf{0.20\pm 0.01}$
DG Adapted IRM $0.92\pm 0.01$ $0.92\pm 0.04$ $0.86\pm 0.04$ $0.32\pm 0.00$ $0.20\pm 0.01$ $0.18\pm 0.01$
Fish $\mathit{0.94\pm 0.02}$ $0.65\pm 0.13$ $0.35\pm 0.11$ $\mathbf{0.35\pm 0.00}$ $0.18\pm 0.00$ $0.15\pm 0.01$
Mixup $0.92\pm 0.01$ $0.89\pm 0.05$ $0.83\pm 0.03$ $0.33\pm 0.01$ $0.09\pm 0.01$ $0.07\pm 0.01$
MMD $0.93\pm 0.02$ $0.92\pm 0.04$ $0.86\pm 0.04$ $0.32\pm 0.00$ $0.19\pm 0.01$ $0.16\pm 0.01$
DeepCoral $0.93\pm 0.01$ $0.92\pm 0.04$ $0.86\pm 0.04$ $0.33\pm 0.01$ $0.19\pm 0.01$ $0.16\pm 0.01$
GroupDRO $0.93\pm 0.01$ $\mathit{0.94\pm 0.03}$ $0.95\pm 0.02$ $0.21\pm 0.00$ $0.13\pm 0.01$ $0.20\pm 0.01$
FL FedProx - $0.90\pm 0.03$ $0.89\pm 0.03$ $0.90\pm 0.02$ $0.25\pm 0.00$ $0.20\pm 0.00$ $0.17\pm 0.00$
Scaffold - $0.90\pm 0.03$ $0.89\pm 0.02$ $0.90\pm 0.02$ $\mathit{0.28\pm 0.00}$ $\mathbf{0.26\pm 0.00}$ $\mathit{0.18\pm 0.01}$
AFL - $0.93\pm 0.03$ $0.92\pm 0.04$ $0.91\pm 0.04$ $0.26\pm 0.01$ $0.16\pm 0.01$ $0.03\pm 0.00$
FDG FedDG - $0.93\pm 0.02$ $\mathit{0.95\pm 0.02}$ $\mathit{0.94\pm 0.02}$ $0.27\pm 0.00$ $0.24\pm 0.00$ $0.17\pm 0.00$
FedADG - $0.94\pm 0.01$ $0.94\pm 0.00$ $0.94\pm 0.01$ $0.26\pm 0.01$ $\mathit{0.25\pm 0.01}$ $0.16\pm 0.00$
FedSR - $0.64\pm 0.16$ $0.55\pm 0.09$ $0.54\pm 0.09$ $0.18\pm 0.01$ $0.14\pm 0.01$ $0.09\pm 0.00$
FedGMA - $0.75\pm 0.05$ $0.73\pm 0.09$ $0.72\pm 0.01$ $0.22\pm 0.00$ $0.15\pm 0.00$ $0.09\pm 0.00$
Table 4: Test accuracy on CelebA and Camelyon17 dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. We report the standard deviation among 3 runs. Please see the Table 9 in the appendix for higher precision report.
CelebA ($D=2$) Camelyon17 ($D=3$)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
B
FedAvg-ERM $0.77\pm 0.04$ $0.61\pm 0.02$ $0.52\pm 0.04$ $0.45\pm 0.02$ $0.90\pm 0.01$ $\mathit{0.95\pm 0.01}$ $\mathit{0.95\pm 0.00}$ $\bf 0.95\pm 0.00$
DG Adapted IRM $\mathbf{0.89\pm 0.01}$ $0.71\pm 0.05$ $0.74\pm 0.01$ $0.91\pm 0.06$ $0.95\pm 0.00$ $0.95\pm 0.00$
Fish $\mathit{0.88\pm 0.01}$ $0.14\pm 0.16$ $0.07\pm 0.05$ $0.92\pm 0.01$ $0.88\pm 0.01$ $0.88\pm 0.01$
Mixup $0.29\pm 0.498$ $0.13\pm 0.22$ $0.13\pm 0.22$ $\mathit{0.94\pm 0.01}$ $\mathbf{0.95\pm 0.00}$ $0.95\pm 0.01$
MMD $0.84\pm 0.05$ $0.81\pm 0.03$ $0.77\pm 0.03$ $0.92\pm 0.03$ $0.95\pm 0.00$ $0.95\pm 0.01$
DeepCoral $0.85\pm 0.04$ $\mathbf{0.81\pm 0.02}$ $\mathbf{0.78\pm 0.02}$ $\mathbf{0.94\pm 0.01}$ $0.95\pm 0.00$ $0.95\pm 0.00$
GroupDRO $0.87\pm 0.03$ $\mathit{0.81\pm 0.03}$ $0.76\pm 0.06$ $0.92\pm 0.03$ $0.95\pm 0.01$ $\mathbf{0.95\pm 0.00}$
FL FedProx - $0.13\pm 0.24$ $0.12\pm 0.25$ $00.00\pm 0$ 0.00 - $0.94\pm 0.00$ $0.94\pm 0.00$ $\mathit{0.94\pm 0.01}$
Scaffold - $0.73\pm 0.01$ $\mathit{0.78\pm 0.02}$ $\bf 0.80\pm 0.02$ $0.94\pm 0.00$ $0.93\pm 0.00$ $0.93\pm 0.00$
AFL - $0.81\pm 0.02$ $0.83\pm 0.01$ $0.83\pm 0.01$ $0.94\pm 0.01$ $0.95\pm 0.01$ $0.93\pm 0.01$
FDG FedDG - $0.56\pm 0.18$ $0.53\pm 0.11$ $0.46\pm 0.20$ $0.86\pm 0.01$ $0.86\pm 0.00$ $0.87\pm 0.02$
FedADG - $0.67\pm 0.00$ $0.67\pm 0.01$ $\mathit{0.60\pm 0.01}$ $0.94\pm 0.00$ $0.93\pm 0.00$ $0.93\pm 0.00$
FedSR - $0.38\pm 0.02$ $0.36\pm 0.04$ $0.38\pm 0.02$ $0.93\pm 0.01$ $0.92\pm 0.01$ $0.93\pm 0.00$
FedGMA - $0.62\pm 0.01$ $0.62\pm 0.02$ $0.49\pm 0.01$ $0.90\pm 0.01$ $0.85\pm 0.01$ $0.82\pm 0.03$
Table 5: Test accuracy on CivilComments and Py150 datasets with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150. We report the standard deviation among 3 runs. Please see the Table 10 in the appendix for higher precision report.
CivilComments ($D=16$) Py150 ($D=5477$)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
B
FedAvg-ERM $0.54\pm 0.00$ $0.36\pm 0.03$ $0.35\pm 0.02$ $0.33\pm 0.01$ $\mathbf{0.68}\pm 0.00$ $\mathbf{0.68}\pm 0.00$ $\mathit{0.65\pm 0.00}$ $\mathit{0.64}\pm 0.00$
DG Adapted IRM $0.64\pm 0.00$ $\mathit{0.59\pm 0.02}$ $0.53\pm 0.04$ $\mathit{0.68}\pm 0.00$ $\mathit{0.66\pm 0.00}$ $\mathbf{0.65\pm 0.00}$ $0.63\pm 0.00$
Fish $\mathbf{0.67}\pm 0.00$ $0.42\pm 0.16$ $0.34\pm 0.17$ $0.66\pm 0.00$ $0.65\pm 0.00$ $0.65\pm 0.00$ $\mathbf{0.64\pm 0.00}$
MMD $\mathit{0.65}\pm 0.00$ $\mathbf{0.63\pm 0.01}$ $\mathbf{0.61\pm 0.01}$ $0.66\pm 0.00$ $0.63\pm 0.00$ $0.63\pm 0.00$ $0.61\pm 0.00$
DeepCoral $0.59\pm 0.00$ $0.52\pm 0.06$ $0.46\pm 0.08$ $0.66\pm 0.00$ $0.65\pm 0.00$ $0.65\pm 0.00$ $0.64\pm 0.00$
GroupDRO $0.64\pm 0.00$ $0.48\pm 0.01$ $\mathit{0.47\pm 0.00}$ $0.51\pm 0.00$ $0.59\pm 0.00$ $0.60\pm 0.00$ $0.61\pm 0.00$
FL FedProx - $0.18\pm 0.03$ $0.17\pm 0.03$ $0.17\pm 0.04$ $0.64\pm 0.00$ $0.63\pm 0.00$ $0.61\pm 0.00$
Scaffold - $0.39\pm 0.02$ $0.38\pm 0.02$ $\mathit{0.33\pm 0.01}$ $0.64\pm 0.00$ $0.64\pm 0.00$ $0.62\pm 0.00$
AFL - $0.55\pm 0.02$ $0.47\pm 0.01$ $\mathbf{0.44\pm 0.02}$ $0.49\pm 0.00$ $0.49\pm 0.00$ $0.47\pm 0.00$
FDG FedSR - $0.36\pm 0.00$ $0.34\pm 0.00$ $0.32\pm 0.00$ $0.53\pm 0.00$ $0.53\pm 0.00$ $0.45\pm 0.00$
FedGMA - $0.21\pm 0.03$ $0.20\pm 0.02$ $0.20\pm 0.02$ $0.62\pm 0.00$ $0.61\pm 0.00$ $0.60\pm 0.00$
[…]